# Exceptional Tannaka groups only arise from cubic threefolds

### v2

@@ Acknowledgements
We thank Victor Gonzalez-Alonso, Manfred Lehn, Luigi Lombardi, Claude Sabbah, Christian Schnell and Mads Villadsen for discussions related to this project. The computer-aided searches in propositions 4.3, LABEL: and 4.5 were carried out by Mads Villadsen. C. L. was supported by the DFG through the research grants Le 3093/3-2, Le 3093/5-1. T. K. was supported by the DFG through the research grant Kr 4663/2-1.
@@ 2.2. Hodge structure on cohomology
[…]
@@ 2.3. Tannaka categories
$*\colon\quad\overline{\textup{Perv}}(A)\times\overline{\textup{Perv}}(A)\;\longrightarrow\;\overline{\textup{Perv}}(A),\quad P_{1}*P_{2}\;:=\;{}^{p}\mathcal{H}^{0}(R\sigma_{*}(P_{1}\boxtimes P_{2}))$
making $\overline{\textup{Perv}}(A)$ a neutral Tannaka category as recalled in [JKLM23, sect. 3.1]. The same constructions work also for the category of Hodge modules: Consider the Serre subcategory
@@
$*\colon\quad\operatorname{\overline{\operatorname{HM}}}(A)\times\operatorname{\overline{\operatorname{HM}}}(A)\;\longrightarrow\;\operatorname{\overline{\operatorname{HM}}}(A),\quad M_{1}*M_{2}\;:=\;{{}_{\scriptscriptstyle{T}}{\sigma}}_{*}^{(0)}(M_{1}\boxtimes M_{2}).$
@@ 2.4. An exact sequence of Tannaka groups
be the full subcategory of objects in $\mathcal{C}$ supported at the origin; it can be seen as a tensor subcategory of the category of bigraded vector spaces via example 2.1, hence its Tannaka group
@@
is a quotient of a torus of rank two; in particular, it is a torus. We then have the following Hodge theoretic analog of the Galois sequence in [JKLM23, th. 4.3]:
@@
$1\longrightarrow G_{\omega}(\operatorname{DR}(\mathcal{C}))\longrightarrow G_{\omega}(\mathcal{C})\longrightarrow G_{\omega}(\mathcal{C}_{\{0\}})\longrightarrow 1$
@@
We have an embedding $\mathcal{C}_{\{0\}}\hookrightarrow\mathcal{C}$ as a full abelian tensor subcategory which is stable under subobjects. By Tannaka duality, any such embedding corresponds to an epimorphism $G_{\omega}(\mathcal{C})\twoheadrightarrow G_{\omega}(\mathcal{C}_{\{0\}})$ [DMOS82, prop. 2.21a]. The exactness of the sequence of Tannaka groups follows by the same argument as in the proof of [JKLM23, th. 4.3]. Indeed, the argument there only needs that for any $M\in\mathcal{C}$ the following two properties hold:
@@
    Any rank one subobject of $\operatorname{DR}(M)$ lies in a semisimple subobject of $\operatorname{DR}(M)$.
@@
Both properties are trivial in the case at hand: The rank one objects of the tensor category $\mathcal{C}$ are the perverse skyscraper sheaves, and the maximal trivial subobject is the direct sum of all skyscraper subsheaves supported at the origin. Hence both claims follow from the fact that for any $M\in\operatorname{HM}(A)$, every simple subobject of the perverse sheaf $\operatorname{DR}(M)$ (or equivalently, of the underlying holonomic $\mathcal{D}$-module) has the property that a direct sum of twists of that subobject underlies a Hodge submodule of $M$ by [SS18, prop. 16.3.1]. ∎
@@
$G_{\omega}(M)$ ::=:= $\operatorname{Im}(G_{\omega}(\mathcal{C})\to\operatorname{GL}(V))$
$G_{\omega}(P)$ ::=:= $\operatorname{Im}(G_{\omega}(\operatorname{DR}(\mathcal{C}))\to\operatorname{GL}(V))$
@@
Applying proposition 2.4 to the tensor category $\mathcal{C}=\langle M\rangle$ generated by $M$, we get an exact sequence
$1\longrightarrow G_{\omega}(P)\longrightarrow G_{\omega}(M)\longrightarrow G_{\omega}(\mathcal{C}_{\{0\}})\longrightarrow 1$
So $G_{\omega}(P)\unlhd G_{\omega}(M)$ is a normal subgroup. Since both are subgroups of $\operatorname{GL}(V)$, this implies the claim about the containment in the normalizer. For the claim about the derived groups of the connected component, recall that $G_{\omega}(\mathcal{C}_{\{0\}})$ is a torus; so the reductive Lie algebras $\operatorname{Lie}G_{\omega}(M)$ and $\operatorname{Lie}G_{\omega}(P)$ have the same maximal semisimple Lie subalgebra. ∎
@@ 2.5. Character twists
which descends to the quotient categories modulo negligibles, where $M\in\operatorname{HM}(A)$ is called negligible if $\operatorname{DR}(M)$ is a negligible perverse sheaf in the sense of section 2.3. The induced functor
@@
As in [KW15, prop. 4.1], twisting by $\mathbb{L}$ gives rise to an equivalence of tensor categories
$\langle M\rangle\;\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}\;\langle M\otimes\mathbb{L}\rangle,\quad N\;\longmapsto\;N\otimes\mathbb{L}$
in $\operatorname{\overline{\operatorname{HM}}}(A)$. This equivalence need not be compatible with the fiber functor $\omega$ on the source and target, but since $\mathbb{C}$ is algebraically closed, any two fiber functors are noncanonically isomorphic; hence the same holds for the Tannaka groups. ∎
@@ 3.1. The Hodge decomposition
As in [JKLM23, sect. 4.3], its image
@@
$\omega\colon\quad\overline{\textup{Perv}}_{0}(A)\;=\;\textup{Perv}_{0}(A)/S_{0}(A)\;\longrightarrow\;\operatorname{Vect}(\mathbb{C}),\quad Q\;\longmapsto\;H^{0}(A,Q).$
This is a tensor functor. In particular, for all perverse sheaves $P_{1},P_{2}\in\overline{\textup{Perv}}(A)$ we have natural isomorphisms
$\vartheta_{P_{1},P_{2}}\colon\quad\omega(P_{1}*P_{2})\;\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}\;\omega(P_{1})\otimes\omega(P_{2})$
@@
$\omega(P)\;=\;\mathcal{H}^{0}(\textup{R}f_{*}(P))\quad\text{for all}\quad P\in\textup{Perv}_{0}(A).$
Now, the complex $\textup{R}f_{\ast}(P)$ has cohomology at most in degree 00 so we identify it with $\omega(P)$. With this in mind, using the Künneth isomorphism and the fact that $f\circ\sigma=f\times f$, we obtain
$\omega(P_{1}*P_{2})$ $\;\simeq\;Rf_{*}(P_{1}*P_{2})$
@@
$\;\simeq\;\omega(P_{1})\otimes\omega(P_{2}).$
These constructions can be upgraded to Hodge modules: Let $\operatorname{HM}_{0}(A)\subset\operatorname{HM}(A)$ be the full subcategory of all Hodge modules $M$ with $\operatorname{DR}(M)\in\textup{Perv}_{0}(A)$. Then $\omega$ lifts to the functor
$\omega^{H}\colon\quad\operatorname{HM}_{0}(A)\;\longrightarrow\;\operatorname{Vect}_{\mathbb{Z}\times\mathbb{Z}}(\mathbb{C}),\quad M\;\longmapsto\;\textup{H}^{0}(A,\operatorname{DR}(M))$
where the bigrading on the values of the functor is the Hodge decomposition from section 2.2:
@@
For $M_{1},M_{2}\in\mathcal{C}$ the isomorphism $\vartheta_{\operatorname{DR}(M_{1}),\operatorname{DR}(M_{2})}$ is compatible with the Hodge decomposition. Hence the functor $\omega^{H}$ is a tensor functor.
@@
$\omega^{H}(M)\;=\;{{}_{\scriptscriptstyle{T}}{f}}^{(0)}_{*}(M)\quad\text{ for all\leavevmode\nobreak\ $M\in\operatorname{HM}_{0}(A)$.}$
As above, the complex of bigraded vector space ${{}_{\scriptscriptstyle{T}}{f}}_{*}(M)$ has cohomology at most in degree 00 so we identify it with $\omega^{H}(M)$. Let $\sigma\colon A\times A\to A$ be the group law. Then we have $f\circ\sigma=f\times f$, hence the Leray spectral sequence [SS18, cor. 12.7.38] gives
@@
$\omega^{H}(M_{1}*M_{2})$ $\;\simeq\;{{}_{\scriptscriptstyle{T}}{f}}_{*}(M_{1}*M_{2})$
@@
($\;\simeq\;{{}_{\scriptscriptstyle{T}}{(f\times f)}}_{*}(M_{1}\boxtimes M_{2})$
@@
$\;\simeq\;\omega^{H}(M_{1})\otimes\omega^{H}(M_{2}),$
@@
The above can be applied to any Hodge module $M\in\operatorname{HM}(A)$ after twisting by a suitable rank one local system. More precisely, by generic vanishing on abelian varieties [KW15, Sch15], we can find a unitary local system $L$ of rank one on $A$ such that
$M\otimes\mathbb{L}\;\in\;\operatorname{HM}_{0}(A).$
Recall that by lemma 2.6, the isomorphism class of the Tannaka group is unaffected by the twist. After the twist, we obtain:
@@
For $M$$\mathbb{L}$ as above, there is a morphism $\lambda\colon\mathbb{G}_{m}^{2}\to G_{\omega}(M\otimes\mathbb{L})$ of algebraic groups such that the Hodge decomposition
$\omega^{H}(M\otimes\mathbb{L})\;=\;\bigoplus_{(p,q)\in\mathbb{Z}^{2}}\textup{H}^{p,q}(M\otimes\mathbb{L})$
@@
The group morphism $\lambda\colon\mathbb{G}_{m}^{2}\to G_{\omega}(\mathcal{C})$ in the above corollary is called the Hodge cocharacter.
The fact that $\lambda$ factors over the Tannaka group will allow us to obtain information about Tannaka groups from information about Hodge numbers.
@@ 3.2. Hodge level of intersection complexes
where the twist is by a unitary local system L of rank one with $\textup{H}^{i}(A,\operatorname{DR}(M)\otimes L)=0$ for all $i\neq 0$; we drop the local system from the notation since it does not affect the Hodge numbers. If $\mathbb{V}=\mathbb{C}_{U}$ is the trivial variation of Hodge structures on the smooth locus $U\subset X$ of an integral subvariety $X\subset A$, we also write
@@
$h^{p}(X)\;\geqslant\;\begin{cases}g-d+1&\text{for\leavevmode\nobreak\ $p\in\{0,d\}$},\\[1.00006pt] 2&\text{for\leavevmode\nobreak\ $p\in\{1,d-1\}$},\\[1.00006pt] 1&\text{for\leavevmode\nobreak\ $2\leqslant p\leqslant d-2$.}\end{cases}$
@@
If $X$ is smooth, then $h^{p}(X)=(-1)^{d-p}\chi(X,\Omega^{p}_{X})$ by [KM23, appendix B], thus $h^{p}(X)=h^{d-p}(X)$ by Serre duality. For the inequalities see corollary A.2. ∎
@@
we may replace $\mathbb{V}$ by the variation of $\mathbb{C}$-Hodge structures $\mathbb{V}\otimes\mathbb{L}$ which has the same level. Hence in what follows we will assume that the unitary local system $\mathbb{L}$ is trivial.
@@
Now let $k=\max\{p\in\mathbb{Z}\mid F^{p}\mathbb{V}\neq 0\}$, so that $F^{k}\mathbb{V}\subset\mathbb{V}$ is the smallest piece in the Hodge filtration. If the variation of Hodge structures has weight $w$, then by Hodge symmetry the Hodge numbers of its fibers can be nonzero at most in bidegrees $(i,w-i)$ with $k\leqslant i\leqslant w-k$. So the level of the Hodge structure on the fibers is
[…]
@@
$F_{p}\mathcal{M}_{U}\;:=\;\omega_{U}\otimes F^{p-d}\mathbb{V}$
which starts in degree $p=k+d$. By the theory of Hodge modules this filtration naturally extends to an increasing filtration on the right $\mathcal{D}_{A}$-module $\mathcal{M}$ underlying the Hodge module $M=\operatorname{IC}_{X}(\mathbb{V})$. Moreover, the lowest piece $F_{p}\mathcal{M}$ of this filtration is still in degree $p=k+d$, see for instance [SY23, sect. 2.4] in the context of Saito’s Hodge modules. Now consider the de Rham complex
@@
which by the above starts in degree $p=k+d$. Therefore, the Hodge filtration on de Rham cohomology
@@
can only start in degree $p=k+d$ or higher. Since $M$ is a pure Hodge module of weight $w+d$ by [SS18, th. 14.6.1], the de Rham cohomology $\textup{H}^{0}(A,\operatorname{DR}(M))$ in degree zero is a pure Hodge structure of weight $w+d$ by th. 14.3.2(2) in loc. cit. So it follows again by Hodge symmetry that its level is
@@
Recall that by $S$-decomposability of Hodge modules [SS18, th. 14.2.17(1)], we have
@@
where $X\subset A$ runs through all integral subvarieties and $M_{X}\subset{{}_{\scriptscriptstyle{T}}f}_{*}^{(0)}(\delta_{Y}^{H})$ denotes the maximal Hodge submodule with strict support $X$. It suffices to prove the lemma when $M=M_{X}$ for some $X\subset A$. Since the strict support decomposition is compatible with the natural functor from Saito’s $\mathbb{Q}$-Hodge modules to complex Hodge modules, the summand $M\subset{{}_{\scriptscriptstyle{T}}f}_{*}^{(0)}(\delta_{Y}^{H})$ then arises from a direct summand in the category of $\mathbb{Q}$-Hodge modules; this will allow us to embed its generic stalk cohomology in the mixed Hodge structure of a singular variety even though the mixed case is not yet written for complex Hodge modules [SS18].
@@
$\ell(\mathbb{V})\;\leqslant\;\ell(\textup{H}^{\bullet}(f^{-1}(Y))\;\leqslant\;\dim f^{-1}(x)\;=\;\dim f^{-1}(X)-d$
where the last equality uses that $x\in X$ is general. Since $M=\operatorname{IC}_{X}(\mathbb{V})$, the previous lemma 3.6 gives $h^{p,q}(M)=0$ for $|p-q|>(\dim f^{-1}(X)-d)+d=\dim f^{-1}(X)$. ∎

### v3

@@ Acknowledgements
We thank Victor Gonzalez-Alonso, Manfred Lehn, Luigi Lombardi, Claude Sabbah, Christian Schnell and Mads Villadsen for discussions related to this project. The computer-aided searches in propositions˜4.3, LABEL: and 4.5 were carried out by Mads Villadsen. We thank the referee for the careful reading and for valuable comments, which helped make the paper more accessible. C. L. was supported by the DFG through the research grants Le 3093/3-2, Le 3093/5-1. T. K. was supported by the DFG through the research grant Kr 4663/2-1.
@@ 2.2. Hodge structure on cohomology
$\operatorname{HM}(X)$$\operatorname{HM}(Y)$$\textup{Perv}(X)$$\textup{Perv}(Y)$$\scriptstyle{{{}_{\scriptscriptstyle{T}}f}_{*}^{(k)}}$$\scriptstyle{\operatorname{DR}}$$\scriptstyle{\operatorname{DR}}$$\scriptstyle{{}^{p}\mathcal{H}^{k}\circ\textup{R}f_{*}}$
@@ 2.3. Tannaka categories
$\colon\quad\overline{\textup{Perv}}(A)\times\overline{\textup{Perv}}(A)\;\longrightarrow\;\overline{\textup{Perv}}(A),\quad P_{1}*P_{2}\;:=\;{}^{p}\mathcal{H}^{0}(R\sigma_{*}(P_{1}\boxtimes P_{2}))$
making $\overline{\textup{Perv}}(A)$ a neutral Tannaka category as recalled in [JKLM25, sect. 3.1]. For any fiber functor $\omega\colon\overline{\textup{Perv}}(A)\to\operatorname{Vect}(\mathbb{C})$, the dimension of objects is the Euler characteristic
(2.1) $\dim_{\mathbb{C}}\,\omega(P)\;=\;\chi(A,P),$
see [KW15b, proof of cor. 4.2]. The same constructions work also for the category of Hodge modules: Consider the Serre subcategory
@@
$\colon\quad\operatorname{\overline{\operatorname{HM}}}(A)\times\operatorname{\overline{\operatorname{HM}}}(A)\;\longrightarrow\;\operatorname{\overline{\operatorname{HM}}}(A),\quad M_{1}*M_{2}\;:=\;{{}_{\scriptscriptstyle{T}}{\sigma}}_{*}^{(0)}(M_{1}\boxtimes M_{2}).$
@@ 2.4. An exact sequence of Tannaka groups
be the full subcategory of objects in $\mathcal{C}$ supported at the origin; it can be seen as a tensor subcategory of the category of bigraded vector spaces via example˜2.1, hence its Tannaka group
@@
is a quotient of a torus of rank two; in particular, it is a torus. We then have the following Hodge theoretic analog of the Galois sequence in [JKLM25, th. 4.3]:
@@
$1\longrightarrow G_{\omega}(\operatorname{DR}(\mathcal{C}))\longrightarrow G_{\omega}(\mathcal{C})\longrightarrow G_{\omega}(\mathcal{C}_{\{0\}})\longrightarrow 1.$
@@
We have an embedding $\mathcal{C}_{\{0\}}\hookrightarrow\mathcal{C}$ as a full abelian tensor subcategory which is stable under subobjects. By Tannaka duality, any such embedding corresponds to an epimorphism $G_{\omega}(\mathcal{C})\twoheadrightarrow G_{\omega}(\mathcal{C}_{\{0\}})$ [DMOS82, prop. 2.21a]. The exactness of the sequence of Tannaka groups then follows by the same argument as in the proof of [JKLM25, th. 4.3]. Indeed, by [DE21, prop. A.13 and lemma A.4(1)] we only need to show that for any $M\in\mathcal{C}$ the following two properties hold:
@@
    Every rank one subobject of $\operatorname{DR}(M)$ is a direct summand in a semisimple object $\operatorname{DR}(N)$ with $N\in\mathcal{C}$.
@@
Both properties are trivial in this case: By [Wei15, sect. 10] or [Krä22, lemma 3.5], the rank one objects of the tensor category $\mathcal{C}$ are precisely those whose underlying perverse sheaf is a skyscraper sheaf of rank one, and the maximal trivial subobject is the direct sum of all skyscraper subsheaves supported at the origin. Hence both claims follow from the fact that for any $M\in\operatorname{HM}(A)$, every simple subobject of the perverse sheaf $\operatorname{DR}(M)$ (or equivalently, of the underlying holonomic $\mathcal{D}$-module) has the property that a direct sum of twists of that subobject underlies a Hodge submodule of $M$ [SS18, prop. 16.3.1]. ∎
@@
$G_{\omega}(M)$ $:=$ $\operatorname{Im}(G_{\omega}(\mathcal{C})\to\operatorname{GL}(V))$
$G_{\omega}(P)$ $:=$ $\operatorname{Im}(G_{\omega}(\operatorname{DR}(\mathcal{C}))\to\operatorname{GL}(V))$
@@
Applying proposition˜2.4 to the tensor category $\mathcal{C}=\langle M\rangle$ generated by $M$, we get an exact sequence
$1\longrightarrow G_{\omega}(P)\longrightarrow G_{\omega}(M)\longrightarrow G_{\omega}(\mathcal{C}_{\{0\}})\longrightarrow 1.$
So $G_{\omega}(P)\unlhd G_{\omega}(M)$ is a normal subgroup. Since both are subgroups of $\operatorname{GL}(V)$, this implies the claim about the containment in the normalizer. For the claim about the derived groups of the connected component, recall that $G_{\omega}(\mathcal{C}_{\{0\}})$ is a torus; so the reductive Lie algebras $\operatorname{Lie}G_{\omega}(M)$ and $\operatorname{Lie}G_{\omega}(P)$ have the same maximal semisimple Lie subalgebra, hence $G_{\omega}(M)^{\circ}$ and $G_{\omega}(P)^{\circ}$ share the same derived subgroup. ∎
@@ 2.5. Character twists
which descends to the quotient categories modulo negligibles, where $M\in\operatorname{HM}(A)$ is called negligible if $\operatorname{DR}(M)$ is a negligible perverse sheaf in the sense of section˜2.3. The induced functor
@@
As in [KW15b, prop. 4.1], twisting by $\mathbb{L}$ gives rise to an equivalence of tensor categories
$\varphi\colon\quad\langle M\rangle\;\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}\;\langle M\otimes\mathbb{L}\rangle,\quad N\;\longmapsto\;N\otimes\mathbb{L}.$
Both the source and the target of this equivalence are tensor subcategories of $\mathcal{C}$, hence $\omega$ restricts to a fiber functor on both of them. The equivalence $\varphi$ need not be compatible with these fiber functors, but over an algebraically closed field of characteristic zero any two fiber functors are noncanonically isomorphic; hence the same holds for the corresponding Tannaka groups. ∎
@@ 3.1. The Hodge decomposition
As in [JKLM25, sect. 4.3], its image
@@
$\omega_{0}\colon\quad\overline{\textup{Perv}}_{0}(A)\;=\;\textup{Perv}_{0}(A)/S_{0}(A)\;\longrightarrow\;\operatorname{Vect}(\mathbb{C}),\quad Q\;\longmapsto\;H^{0}(A,Q).$
This is a tensor functor. In particular, for all perverse sheaves $P_{1},P_{2}\in\overline{\textup{Perv}}_{0}(A)$ we have natural isomorphisms
$\vartheta_{P_{1},P_{2}}\colon\quad\omega_{0}(P_{1}*P_{2})\;\stackrel{{\scriptstyle\sim}}{{\longrightarrow}}\;\omega_{0}(P_{1})\otimes\omega_{0}(P_{2})$
@@
$\omega_{0}(P)\;=\;\mathcal{H}^{0}(\textup{R}f_{*}(P))\quad\text{for all}\quad P\in\textup{Perv}_{0}(A).$
Now, the complex $\textup{R}f_{\ast}(P)$ has cohomology at most in degree 0 so we identify it with $\omega(P)$. With this in mind, using the Künneth isomorphism and the fact that $f\circ\sigma=f\times f$, we obtain
$\omega_{0}(P_{1}*P_{2})$ $\;\simeq\;Rf_{*}(P_{1}*P_{2})$
@@
$\;\simeq\;\omega_{0}(P_{1})\otimes\omega_{0}(P_{2}).$
These constructions can be upgraded to Hodge modules: Let $\operatorname{HM}_{0}(A)\subset\operatorname{HM}(A)$ be the full subcategory of all Hodge modules $M$ with $\operatorname{DR}(M)\in\textup{Perv}_{0}(A)$, and denote by $\operatorname{\overline{\operatorname{HM}}}_{0}(A)\subset\operatorname{\overline{\operatorname{HM}}}(A)$ its image in the Tannaka category from section˜2.3. Then $\omega_{0}$ induces the functor
$\omega_{0}^{H}\colon\quad\operatorname{\overline{\operatorname{HM}}}_{0}(A)\;\longrightarrow\;\operatorname{Vect}_{\mathbb{Z}\times\mathbb{Z}}(\mathbb{C}),\quad M\;\longmapsto\;\textup{H}^{0}(A,\operatorname{DR}(M))$
where the bigrading on the values of the functor is the Hodge decomposition from section˜2.2:
@@
For $M_{1},M_{2}\in\operatorname{\overline{\operatorname{HM}}}_{0}(A)$, the isomorphism $\vartheta_{\operatorname{DR}(M_{1}),\operatorname{DR}(M_{2})}$ is compatible with the Hodge decomposition. Hence the functor $\omega_{0}^{H}$ is a tensor functor.
@@
$\omega_{0}^{H}(M)\;=\;{{}_{\scriptscriptstyle{T}}{f}}^{(0)}_{*}(M)\quad\text{ for all\penalty 10000\ $M\in\operatorname{\overline{\operatorname{HM}}}_{0}(A)$.}$
As above, the complex of bigraded vector space ${{}_{\scriptscriptstyle{T}}{f}}_{*}(M)$ has cohomology at most in degree 0 so we identify it with $\omega_{0}^{H}(M)$. Let $\sigma\colon A\times A\to A$ be the group law. Then we have $f\circ\sigma=f\times f$, hence the Leray spectral sequence [SS18, cor. 12.7.38] gives
@@
$\omega_{0}^{H}(M_{1}*M_{2})$ $\;\simeq\;{{}_{\scriptscriptstyle{T}}{f}}_{*}(M_{1}*M_{2})$
@@
$\;\simeq\;{{}_{\scriptscriptstyle{T}}{(f\times f)}}_{*}(M_{1}\boxtimes M_{2})$
@@
$\;\simeq\;\omega_{0}^{H}(M_{1})\otimes\omega_{0}^{H}(M_{2}),$
@@
The above can be applied to any Hodge module $M\in\operatorname{\overline{\operatorname{HM}}}(A)$ after twisting by a suitable rank one local system. More precisely, by generic vanishing on abelian varieties [KW15b, Sch15] we can find a unitary local system $L$ of rank one on $A$ such that
$M\otimes\mathbb{L}\;\in\;\operatorname{\overline{\operatorname{HM}}}_{0}(A).$
Fixing $M$ and $\mathbb{L}$, we obtain a tensor functor
$\omega^{H}\colon\quad\langle M\rangle\;\longrightarrow\;\operatorname{Vect}_{\mathbb{Z}\times\mathbb{Z}}(\mathbb{C}),\quad N\;\longmapsto\;\textup{H}^{0}(A,N\otimes\mathbb{L})$
by precomposing $\omega_{0}^{H}$ with the tensor functor $N\mapsto N\otimes\mathbb{L}$, see [KW15b, prop. 4.1].
@@
There is a morphism $\lambda\colon\mathbb{G}_{m}^{2}\to G_{\omega}(M)$ of algebraic groups such that the Hodge decomposition
@@
The morphism $\lambda\colon\mathbb{G}_{m}^{2}\to G_{\omega}(M)$ in the above corollary is called the Hodge cocharacter.
The fact that $\lambda$ factors over the Tannaka group will allow us to find information about Tannaka groups from information about Hodge numbers.
@@ 3.2. Hodge level of intersection complexes
where the twist is by a unitary local system L of rank one with $\textup{H}^{i}(A,\operatorname{DR}(M)\otimes L)=0$ for all $i\neq 0$; we drop the local system from the notation since it does not affect the Hodge numbers. If $\mathbb{V}=\mathbb{C}_{U}$ is the trivial variation of Hodge structures on the smooth locus $U\subset X$ of an integral subvariety $X\subset A$ of dimension $d$, we also write
@@
$h^{p}(X)\;\geqslant\;\begin{cases}g-d+1&\text{for\penalty 10000\ $p\in\{0,d\}$},\\[1.00006pt] 2&\text{for\penalty 10000\ $p\in\{1,d-1\}$},\\[1.00006pt] 1&\text{for\penalty 10000\ $2\leqslant p\leqslant d-2$.}\end{cases}$
@@
If $X$ is smooth, then $h^{p}(X)=(-1)^{d-p}\chi(X,\Omega^{p}_{X})$ by [KM23, appendix B], thus $h^{p}(X)=h^{d-p}(X)$ by Serre duality. For the inequalities see corollary˜A.2. ∎
@@
we may replace $\mathbb{V}$ by the variation of $\mathbb{C}$-Hodge structures $\mathbb{V}\otimes\mathbb{L}$ which has the same level. Hence in what follows we will assume that $\mathbb{L}$ is the trivial local system.
@@
Now let $k=\max\{p\in\mathbb{Z}\mid F^{p}\mathbb{V}\neq 0\}$, so that $F^{k}\mathbb{V}\subset\mathbb{V}$ is the smallest piece in the Hodge filtration. If the variation of Hodge structures has weight $w$, then by Hodge symmetry the Hodge numbers of its fibers can be nonzero at most in bidegrees (i,wi)(i,w-i) with $w-k\leqslant i\leqslant k$. So the level of the Hodge structure on the fibers is
$\ell(\mathbb{V})\;=\;2k-w.$
@@
$F_{p}\mathcal{M}_{U}\;:=\;\omega_{U}\otimes F^{-p-d}\mathbb{V}$
which starts in degree p=kdp=-k-d. By the theory of Hodge modules this filtration naturally extends to an increasing filtration on the right $\mathcal{D}_{A}$-module $\mathcal{M}$ underlying the Hodge module $M=\operatorname{IC}_{X}(\mathbb{V})$. Moreover, the lowest piece $F_{p}\mathcal{M}$ of this filtration is still in degree p=kdp=-k-d, see for instance [SY23, sect. 2.4] in the context of Saito’s Hodge modules. Now consider the de Rham complex
@@
which by the above starts in degree p=kdp=-k-d. Hence on de Rham cohomology the filtration
@@
can only start in degree p=kdp=-k-d or higher. Since $M$ is a pure Hodge module of weight $w+d$ by [SS18, th. 14.6.1], the de Rham cohomology $\textup{H}^{0}(A,\operatorname{DR}(M))$ in degree zero is a pure Hodge structure of weight $w+d$ by th. 14.3.2(2) in loc. cit. Its classical decreasing Hodge filtration goes at most up to degree p=k+d-p=k+d by the above. So again by Hodge symmetry its level is
@@
Recall that by $S$-decomposability of Hodge modules [SS18, th. 14.2.19(1)], we have
@@
where $X\subset A$ runs through all integral subvarieties and $M_{X}\subset{{}_{\scriptscriptstyle{T}}f}_{*}^{(0)}(\delta_{Y}^{H})$ denotes the maximal Hodge submodule with strict support $X$. It thus suffices to prove the lemma when $M=M_{X}$ for some $X\subset A$. Note that the above strict support decomposition is compatible with the natural functor from Saito’s $\mathbb{Q}$-Hodge modules to complex Hodge modules, since under the faithful de Rham functor it corresponds to the strict support decomposition of the underlying perverse sheaves. Thus, the summand $M\subset{{}_{\scriptscriptstyle{T}}f}_{*}^{(0)}(\delta_{Y}^{H})$ arises from a direct summand in the category of $\mathbb{Q}$-Hodge modules; this will allow us to embed its generic stalk cohomology in the mixed Hodge structure of a singular variety even though the mixed case is not yet written for complex Hodge modules [SS18].
@@
$\ell(\mathbb{V})\;\leqslant\;\ell(\textup{H}^{\bullet}(f^{-1}(x))\;\leqslant\;\dim f^{-1}(x)\;=\;\dim f^{-1}(X)-d,$
where the first inequality comes from the inclusion $V_{x}\hookrightarrow\textup{H}^{\bullet}(f^{-1}(x))$ of Hodge structures and the last equality uses that $x\in X$ is general. Since $M=\operatorname{IC}_{X}(\mathbb{V})$, lemma˜3.6 gives $h^{p,q}(M)=0$ for $|p-q|>(\dim f^{-1}(X)-d)+d=\dim f^{-1}(X)$. ∎